OpenStudy (anonymous):

cosxtanx - 2cos^2x = -1

OpenStudy (anonymous):

does it help to know that \(\cos(x)\tan(x)=\sin(x)\)?

OpenStudy (tkhunny):

\(\cos^{2}(x) = 1 - \sin^{2}(x)\) Sines are wild!